# 6D-Diff: A Keypoint Diffusion Framework for 6D Object Pose Estimation

### 02 Jan 2024 v2

@@ Abstract
Estimating the 6D object pose from a single RGB image often involves noise and indeterminacy due to challenges such as occlusions and cluttered backgrounds. Meanwhile, diffusion models have shown appealing performance in generating high-quality images from random noise with high indeterminacy through step-by-step denoising. Inspired by their denoising capability, we propose a novel diffusion-based framework (6D-Diff) to handle the noise and indeterminacy in object pose estimation for better performance. In our framework, to establish accurate 2D-3D correspondence, we formulate 2D keypoints detection as a reverse diffusion (denoising) process. To facilitate such a denoising process, we design a Mixture-of-Cauchy-based forward diffusion process and condition the reverse process on the object features. Extensive experiments on the LM-O and YCB-V datasets demonstrate the effectiveness of our framework.
@@ 1 Introduction
6D object pose estimation aims to estimate the 6D pose of an object including its location and orientation, which has a wide range of applications, such as augmented reality [37, 45], robotic manipulation [43, 3], and automatic driving [60]. Recently, various methods [42, 62, 51, 4, 25, 21, 59, 5, 18] have been proposed to conduct RGB-based 6D object pose estimation since RGB images are easy to obtain. Despite the increased efforts, a variety of challenges persist in RGB-based 6D object pose estimation, including occlusions, cluttered backgrounds, and changeable environment [61, 42, 8, 58, 38]. These challenges can introduce significant noise and indeterminacy into the pose estimation process, leading to error-prone predictions [42, 8, 38].
@@
In the forward process, we aim to generate supervision signals that will be used to optimize the diffusion model during the reverse process. Specifically, given a set of pre-selected 3D keypoints, we first acquire ground-truth coordinates of their corresponding 2D keypoints using the ground-truth object pose. Then these determinate ground-truth 2D coordinates are gradually diffused towards the indeterminate distribution modeled from the intermediate representation, and the distributions generated along the way will be used as supervision signals. Note that, as the distribution modeled from the intermediate representation can be complex and irregular, it is difficult to characterize such a distribution via the Gaussian distribution. This means that simply applying diffusion models in most existing generation works [17, 50, 7], which start denoising from the random Gaussian noise, can introduce potentially large error. To tackle this challenge, we draw inspiration from the fact that the Mixture of Cauchy (MoC) model can characterize complex and intractable distributions in an effective and robust manner [24]. Thus we propose to model the intermediate representation using a MoC distribution instead of simply treating it as a random Gaussian noise. In this way, we gradually diffuse the determinate distribution (ground truth) of keypoints coordinates towards the modeled MoC distribution during the forward process.
@@
Figure 2: Above we show two examples of keypoint heatmaps, which serve as the intermediate representation [5, 42, 4] in our framework. The red dots indicate the ground-truth locations of the keypoints. As shown above, due to occlusions and cluttered backgrounds, the keypoint heatmaps are noisy, which reflects the noise and indeterminacy during the keypoints detection process.
@@
Correspondingly, in the reverse process, starting from the MoC distribution modeled in the forward process, we aim to learn to recover the ground-truth keypoints coordinates. To achieve this, we leverage the distributions generated step-by-step during the forward process as the supervision signals to train the diffusion model to learn the reverse process. In this way, the diffusion model can learn to convert the indeterminate MoC distribution of keypoints coordinates into a determinate one smoothly and effectively. After the reverse process, the 2D keypoints coordinates obtained from the final determinate distribution are used to compute the 6D object pose with the pre-selected 3D keypoints. Moreover, we further facilitate the model learning of such a reverse process by injecting object features as context information.
@@
Our work makes three main contributions. 1) We propose a novel 6D-Diff framework, in which we formulate keypoints detection for 6D object pose estimation as a reverse diffusion process to effectively eliminate the noise and indeterminacy in object pose estimation. 2) To take advantage of the intermediate representation that encodes useful prior distribution knowledge for handling this task, we propose a novel MoC-based diffusion process. Besides, we facilitate the model learning by utilizing object features. 3) Our framework achieves state-of-the-art performance on the evaluated benchmarks.
@@ 2 Related Work
RGB-based 6D Object Pose Estimation has received a lot of attention [61, 41, 42, 44, 54, 51, 4, 22, 31, 36, 52, 64, 62, 15, 30, 34, 63, 12, 14, 13]. Some works [61, 25, 21, 59] proposed to directly regress object poses. However, the non-linearity of the rotation space makes direct regression of object poses difficult [30]. Compared to this type of direct methods, correspondence-based methods [5, 18, 41, 42, 44, 54, 51] often demonstrate better performance, which estimate 6D object poses via learning 2D-3D correspondences between the observed image and the object 3D model.
@@
Diffusion Models [17, 50, 7, 48] are originally introduced for image synthesis. Showing appealing generation capabilities, diffusion models have also been explored in various other tasks [39, 11, 35, 10, 56, 28, 19], such as image editing [39] and image impainting [35]. Here we explore a new framework that tackles object pose estimation with a diffusion model. Different from previous generation works [7, 39, 35] that start denoising from random noise, to aid the denoising process for 6D object pose estimation, we design a novel MoC-based diffusion mechanism that enables the diffusion model to start denoising from a distribution containing useful prior distribution knowledge regarding the object pose. Moreover, we condition the denoising process on the object features, to further guide the diffusion model to obtain accurate predictions.
@@ 3 Method
To handle the noise and indeterminacy in RGB-based 6D object pose estimation, from a novel perspective of distribution transformation with progressive denoising, we propose a framework (6D-Diff) that represents a new brand of diffusion-based solution for 6D object pose estimation. Below we first revisit diffusion models in Sec. 3.1. Then we discuss our proposed framework in Sec. 3.2, and introduce its training and testing scheme in Sec. 3.3. We finally detail the model architecture in Sec. 3.4.
@@ 3.1 Revisiting Diffusion Models
where $\{\beta_{k}\in(0,1)\}^{K}_{k=1}$ denotes a set of fixed variance controllers that control the scale of the injected noise at different steps. According to Eq. 1, we can derive $q(d_{k}|d_{0})$ in closed form as:
@@
where $\alpha_{k}=1-\beta_{k}$ and $\overline{\alpha}_{k}=\prod^{k}_{s=1}\alpha_{s}$. Based on Eq. 2, $d_{k}$ can be further expressed as:
@@
where $\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I})$. From Eq. 3, we can observe that when the number of diffusion steps $K$ is sufficiently large and $\overline{\alpha}_{K}$ correspondingly decreases to nearly zero, the distribution of $d_{K}$ is approximately a standard Gaussian distribution, i.e., $d_{K}\sim\mathcal{N}(\textbf{0},\textbf{I})$. This means $d_{0}$ is gradually corrupted into Gaussian noise, which conforms to the non-equilibrium thermodynamics phenomenon of the diffusion process [48].
@@
Figure 3: Illustration of our framework. During testing, given an input image, we first crop the Region of Interest (ROI) from the image through an object detector. After that, we feed the cropped ROI to the keypoints distribution initializer to obtain the heatmaps that can provide useful distribution priors about keypoints, to initialize $D_{K}$. Meanwhile, we can obtain object features $f_{app}$. Next, we pass $f_{app}$ into the encoder, and the output of the encoder will serve as conditional information to aid the reverse process in the decoder. We sample $M$ sets of 2D keypoints coordinates from $D_{K}$, and feed these $M$ sets of coordinates into the decoder to perform the reverse process iteratively together with the step embedding $f^{k}_{D}$. At the final reverse step ($K$-th step), we average $\{d^{i}_{0}\}^{M}_{i=1}$ as the final keypoints coordinates prediction $d_{0}$, and use $d_{0}$ to compute the 6D pose with the pre-selected 3D keypoints via a PnP solver.
@@ 3.2 Proposed Framework
When detecting 2D keypoints, factors like occlusions (including self-occlusions) and cluttered backgrounds can bring noise and indeterminacy into this process, and affect the accuracy of detection results [42, 20]. To handle this problem, inspired by that diffusion models can iteratively reduce indeterminacy and noise in the initial distribution (e.g., standard Gaussian distribution) to generate determinate and high-quality samples of the desired distribution [11, 10], we formulate keypoints detection as generating a determinate distribution of keypoints coordinates ($D_{0}$) from an indeterminate initial distribution ($D_{K}$) via a diffusion model.
@@
However, it is non-trivial to design such a forward process, as the initialized distribution $D_{K}$ is based on extracted heatmaps, and thus $D_{K}$ can be complex and irregular, as shown in Fig. 4. Hence modeling $D_{K}$ as a Gaussian distribution can result in potentially large error. To handle this challenge, motivated by that the Mixture of Cauchy (MoC) model can effectively and reliably characterize complex and intractable distributions [24], we leverage MoC to characterize $D_{K}$. Based on the characterized distribution, we can then perform a corresponding MoC-based forward process.
@@
Specifically, we denote the number of Cauchy kernels in the MoC distribution as $U$, and use the Expectation-Maximum-type ($EM$) algorithm [24, 53] to optimize the MoC parameters $\eta^{MoC}$ to characterize the distribution $D_{K}$ as:
$\eta_{*}^{MoC}=EM\Big{(}\prod^{V}_{v=1}\sum^{U}_{u=1}\pi_{u}Cauchy(d^{v}_{K}|\mu_{u},\gamma_{u})\Big{)}$ (4)
@@
where $\{d^{v}_{K}\}^{V}_{v=1}$ denotes $V$ sets of keypoints coordinates sampled from the distribution $D_{K}$. Note each set of keypoints coordinates $d^{v}_{K}$ contains all the $N$ keypoints coordinates (i.e., $d^{v}_{K}\in\mathbb{R}^{N\times 2}$). $\pi_{u}$ denotes the weight of the $u$-th Cauchy kernel ($\sum^{U}_{u=1}\pi_{u}$ = 1), and $\eta^{MoC}=\{\mu_{1},\gamma_{1},...,\mu_{U},\gamma_{U}\}$ denotes the MoC parameters in which $\mu_{u}$ and $\gamma_{u}$ are the location and scale of the $u$-th Cauchy kernel. Via the above optimization, we can use the optimized parameters $\eta_{*}^{MoC}$ to model $D_{K}$ as the characterized distribution ($\hat{D}_{K}$). Given $\hat{D}_{K}$, we aim to conduct the forward process from the ground-truth keypoints coordinates distribution $D_{0}$, so that after $K$ steps of forward diffusion, the generated distribution reaches $\hat{D}_{K}$. To this end, we modify Eq. 3 as follows:
$\hat{d}_{k}=\sqrt{\overline{\alpha}_{k}}d_{0}+(1-\sqrt{\overline{\alpha}_{k}})\mu^{MoC}+\sqrt{1-\overline{\alpha}_{k}}\epsilon^{MoC}$ (5)
@@
where $\hat{d}_{k}\in\mathbb{R}^{N\times 2}$ represents a sample (i.e., a set of $N$ keypoints coordinates) from the generated distribution $\hat{D}_{k}$, $\mu^{MoC}=\sum^{U}_{u=1}\mathds{1}_{u}\mu_{u}$, and $\epsilon^{MoC}\sim Cauchy(\textbf{0},\sum^{U}_{u=1}(\mathds{1}_{u}\gamma_{u}))$. Note that $\mathds{1}_{u}$ is a zero-one indicator and $\sum^{U}_{u=1}\mathds{1}_{u}=1$ and $Prob(\mathds{1}_{u}=1)=\pi_{u}$.
@@
From Eq. 5, we can observe that when $K$ is sufficiently large and $\overline{\alpha}_{K}$ correspondingly decreases to nearly zero, the distribution of $\hat{d}_{K}$ reaches the MoC distribution, i.e., $\hat{d}_{K}=\mu^{MoC}+\epsilon^{MoC}\sim Cauchy(\sum^{U}_{u=1}(\mathds{1}_{u}\mu_{u}),\sum^{U}_{u=1}(\mathds{1}_{u}\gamma_{u}))$. After the above MoC-based forward process, we can use the generated $\{\hat{D}_{k}\}^{K-1}_{k=1}$ as supervision signals to train the diffusion model $M_{diff}$ to learn the reverse process. More details about Eq. 5 can be found in Supplementary. Note that such a forward process is only conducted to generate supervision signals for training the diffusion model, while we only need to conduct the reverse process during testing.
@@
However, it can still be difficult to generate $D_{0}$ by directly performing the reverse process from $\hat{D}_{K}$, because the object appearance features are lacking in $\hat{D}_{K}$. Such features can help constrain the model reverse process based on the input image to get accurate predictions. Thus we further leverage the object appearance features as context to guide $M_{diff}$ in the reverse process. Specifically, we seek help from the keypoints distribution initializer to formulate the object features $f_{app}$ and feed $f_{app}$ into the diffusion model, as shown in Fig. 3.
Our reverse process aims to generate a determinate distribution $D_{0}$ from the indeterminate distribution $\hat{D}_{K}$ (during training) or $D_{K}$ (during testing). Below we describe the reverse process during testing. We first obtain $f_{app}$ from the input image. Then to help the diffusion model to learn to perform denoising at each reverse step, following [17, 50], we generate the unique step embedding $f^{k}_{D}$ to inject the step number ($k$) information into the model. In this way, given a set of noisy keypoints coordinates $d_{k}\in\mathbb{R}^{N\times 2}$ drawn from ${D}_{k}$ at the $k^{th}$ step, we use diffusion model $M_{diff}$, conditioned on the step embedding $f^{k}_{D}$ and the object features $f_{app}$, to recover ${d}_{k-1}$ from ${d}_{k}$ as:
${d}_{k-1}=M_{diff}({d}_{k},f_{app},f^{k}_{D})$ (6)

### 22 Mar 2024 v3

@@ Abstract
Estimating the 6D object pose from a single RGB image often involves noise and indeterminacy due to challenges such as occlusions and cluttered backgrounds. Meanwhile, diffusion models have shown appealing performance in generating high-quality images from random noise with high indeterminacy through step-by-step denoising. Inspired by their denoising capability, we propose a novel diffusion-based framework (6D-Diff) to handle the noise and indeterminacy in object pose estimation for better performance. In our framework, to establish accurate 2D-3D correspondence, we formulate 2D keypoints detection as a reverse diffusion (denoising) process. To facilitate such a denoising process, we design a Mixture-of-Cauchy-based forward diffusion process and condition the reverse process on the object appearance features. Extensive experiments on the LM-O and YCB-V datasets demonstrate the effectiveness of our framework.
† Equal contribution;  ‡ Corresponding author
@@ 1 Introduction
6D object pose estimation aims to estimate the 6D pose of an object including its location and orientation, which has a wide range of applications, such as augmented reality [39, 47], robotic manipulation [45, 3], and automatic driving [62]. Recently, various methods [44, 64, 53, 4, 27, 22, 61, 5, 19] have been proposed to conduct RGB-based 6D object pose estimation since RGB images are easy to obtain. Despite the increased efforts, a variety of challenges persist in RGB-based 6D object pose estimation, including occlusions, cluttered backgrounds, and changeable environments [63, 44, 8, 60, 40]. These challenges can introduce significant noise and indeterminacy into the pose estimation process, leading to error-prone predictions [44, 8, 40].
@@
In the forward process, we aim to generate supervision signals that will be used to optimize the diffusion model during the reverse process. Specifically, given a set of pre-selected 3D keypoints, we first acquire ground-truth coordinates of their corresponding 2D keypoints using the ground-truth object pose. Then these determinate ground-truth 2D coordinates are gradually diffused towards the indeterminate distribution modeled from the intermediate representation, and the distributions generated along the way will be used as supervision signals. Note that, as the distribution modeled from the intermediate representation can be complex and irregular, it is difficult to characterize such a distribution via the Gaussian distribution. This means that simply applying diffusion models in most existing generation works [18, 52, 7], which start denoising from the random Gaussian noise, can introduce potentially large errors. To tackle this challenge, we draw inspiration from the fact that the Mixture of Cauchy (MoC) model can effectively characterize complex and intractable distributions. Moreover, the MoC model is robust to potential outliers in the distribution to be characterized [26]. Thus we propose to model the intermediate representation using a MoC distribution instead of simply treating it as a random Gaussian noise. In this way, we gradually diffuse the determinate distribution (ground truth) of keypoints coordinates towards the modeled MoC distribution during the forward process.
@@
Figure 2: Above we show two examples of keypoint heatmaps, which serve as the intermediate representation [5, 44, 4] in our framework. The red dots indicate the ground-truth locations of the keypoints. In the example (a), the target object is the pink cat, which is heavily occluded in the image and is shown in a different pose compared to the 3D model. As shown above, due to occlusions and cluttered backgrounds, the keypoint heatmaps are noisy, which reflects the noise and indeterminacy during the keypoints detection process.
@@
Correspondingly, in the reverse process, starting from the MoC distribution modeled in the forward process, we aim to learn to recover the ground-truth keypoints coordinates. To achieve this, we leverage the distributions generated step-by-step during the forward process as the supervision signals to train the diffusion model to learn the reverse process. In this way, the diffusion model can learn to convert the indeterminate MoC distribution of keypoints coordinates into a determinate one smoothly and effectively. After the reverse process, the 2D keypoints coordinates obtained from the final determinate distribution are used to compute the 6D object pose with the pre-selected 3D keypoints. Moreover, we further facilitate the model learning of such a reverse process by injecting object appearance features as context information.
@@
Our work makes the following contributions. 1) We propose a novel 6D-Diff framework, in which we formulate keypoints detection for 6D object pose estimation as a reverse diffusion process to effectively eliminate the noise and indeterminacy in object pose estimation. 2) To take advantage of the intermediate representation that encodes useful prior distribution knowledge for handling this task, we propose a novel MoC-based diffusion process. Besides, we facilitate the model learning by utilizing object features.
@@ 2 Related Work
RGB-based 6D Object Pose Estimation has received a lot of attention [63, 43, 44, 46, 56, 53, 4, 23, 33, 38, 54, 67, 64, 16, 32, 36, 65, 13, 15, 14, 66]. Some works [63, 27, 22, 61] proposed to directly regress object poses. However, the non-linearity of the rotation space makes direct regression of object poses difficult [32]. Compared to this type of direct methods, correspondence-based methods [5, 19, 43, 44, 46, 56, 53] often demonstrate better performance, which estimate 6D object poses via learning 2D-3D correspondences between the observed image and the object 3D model.
@@
Diffusion Models [18, 52, 7, 50, 9] are originally introduced for image synthesis. Showing appealing generation capabilities, diffusion models have also been explored in various other tasks [41, 12, 37, 11, 25, 58, 30, 20], such as image editing [41] and image impainting [37]. Here we explore a new framework that tackles object pose estimation with a diffusion model. Different from previous generation works [7, 41, 37] that start denoising from random noise, to aid the denoising process for 6D object pose estimation, we design a novel MoC-based diffusion mechanism that enables the diffusion model to start denoising from a distribution containing useful prior distribution knowledge regarding the object pose. Moreover, we condition the denoising process on the object appearance features, to further guide the diffusion model to obtain accurate predictions.
@@ 3 Method
To handle the noise and indeterminacy in RGB-based 6D object pose estimation, inspired by [11], from a novel perspective of distribution transformation with progressive denoising, we propose a framework (6D-Diff) that represents a new brand of diffusion-based solution for 6D object pose estimation. Below we first revisit diffusion models in Sec. 3.1. Then we discuss our proposed framework in Sec. 3.2, and introduce its training and testing scheme in Sec. 3.3. We finally detail the model architecture in Sec. 3.4.
@@ 3.1 Revisiting Diffusion Models
where $\{\beta_{k}\in(0,1)\}^{K}_{k=1}$ denotes a set of fixed variance controllers that control the scale of the injected noise at different steps. According to Eq. (1), we can derive $q(d_{k}|d_{0})$ in closed form as:
@@
where $\alpha_{k}=1-\beta_{k}$ and $\overline{\alpha}_{k}=\prod^{k}_{s=1}\alpha_{s}$. Based on Eq. (2), $d_{k}$ can be further expressed as:
@@
where $\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I})$. From Eq. (3), we can observe that when the number of diffusion steps $K$ is sufficiently large and $\overline{\alpha}_{K}$ correspondingly decreases to nearly zero, the distribution of $d_{K}$ is approximately a standard Gaussian distribution, i.e., $d_{K}\sim\mathcal{N}(\textbf{0},\textbf{I})$. This means $d_{0}$ is gradually corrupted into Gaussian noise, which conforms to the non-equilibrium thermodynamics phenomenon of the diffusion process [50].
@@
Figure 3: Illustration of our framework. During testing, given an input image, we first crop the Region of Interest (ROI) from the image through an object detector. After that, we feed the cropped ROI to the keypoints distribution initializer to obtain the heatmaps that can provide useful distribution priors about keypoints, to initialize $D_{K}$. Meanwhile, we can obtain object appearance features $f_{\text{app}}$. Next, we pass $f_{\text{app}}$ into the encoder, and the output of the encoder will serve as conditional information to aid the reverse process in the decoder. We sample $M$ sets of 2D keypoints coordinates from $D_{K}$, and feed these $M$ sets of coordinates into the decoder to perform the reverse process iteratively together with the step embedding $f^{k}_{D}$. At the final reverse step ($K$-th step), we average $\{d^{i}_{0}\}^{M}_{i=1}$ as the final keypoints coordinates prediction $d_{0}$, and use $d_{0}$ to compute the 6D pose with the pre-selected 3D keypoints via a PnP solver.
@@ 3.2 Proposed Framework
When detecting 2D keypoints, factors like occlusions and cluttered backgrounds can bring noise and indeterminacy into this process, and affect the accuracy of detection results [44, 21]. To handle this problem, inspired by that diffusion models can iteratively reduce indeterminacy and noise in the initial distribution (e.g., standard Gaussian distribution) to generate determinate and high-quality samples of the desired distribution [12, 11], we formulate keypoints detection as generating a determinate distribution of keypoints coordinates ($D_{0}$) from an indeterminate initial distribution ($D_{K}$) via a diffusion model.
@@
However, it is non-trivial to design such a forward process, as the initialized distribution $D_{K}$ is based on extracted heatmaps, and thus $D_{K}$ can be complex and irregular, as shown in Fig. 4. Hence modeling $D_{K}$ as a Gaussian distribution can result in potentially large errors. To handle this challenge, motivated by that the Mixture of Cauchy (MoC) model can effectively and reliably characterize complex and intractable distributions [26], we leverage MoC to characterize $D_{K}$. Based on the characterized distribution, we can then perform a corresponding MoC-based forward process.
@@
Specifically, we denote the number of Cauchy kernels in the MoC distribution as $U$, and use the Expectation-Maximum-type (EM) algorithm [26, 55] to optimize the MoC parameters $\eta^{\text{MoC}}$ to characterize the distribution $D_{K}$ as:
$\eta_{*}^{\text{MoC}}=\text{EM}\Big{(}\prod^{V}_{v=1}\sum^{U}_{u=1}\pi_{u}\text{Cauchy}(d^{v}_{K}|\mu_{u},\gamma_{u})\Big{)}$ (4)
@@
where $\{d^{v}_{K}\}^{V}_{v=1}$ denotes $V$ sets of keypoints coordinates sampled from the distribution $D_{K}$. Note each set of keypoints coordinates $d^{v}_{K}$ contains all the $N$ keypoints coordinates (i.e., $d^{v}_{K}\in\mathbb{R}^{N\times 2}$). $\pi_{u}$ denotes the weight of the $u$-th Cauchy kernel ($\sum^{U}_{u=1}\pi_{u}$ = 1), and $\eta^{\text{MoC}}=\{\mu_{1},\gamma_{1},...,\mu_{U},\gamma_{U}\}$ denotes the MoC parameters in which $\mu_{u}$ and $\gamma_{u}$ are the location and scale of the $u$-th Cauchy kernel. Via the above optimization, we can use the optimized parameters $\eta_{*}^{\text{MoC}}$ to model $D_{K}$ as the characterized distribution ($\hat{D}_{K}$). Given $\hat{D}_{K}$, we aim to conduct the forward process from the ground-truth keypoints coordinates distribution $D_{0}$, so that after $K$ steps of forward diffusion, the generated distribution reaches $\hat{D}_{K}$. To this end, we modify Eq. (3) as follows:
$\hat{d}_{k}=\sqrt{\overline{\alpha}_{k}}d_{0}+(1-\sqrt{\overline{\alpha}_{k}})\mu^{\text{MoC}}+\sqrt{1-\overline{\alpha}_{k}}\epsilon^{\text{MoC}}$ (5)
@@
where $\hat{d}_{k}\in\mathbb{R}^{N\times 2}$ represents a sample (i.e., a set of $N$ keypoints coordinates) from the generated distribution $\hat{D}_{k}$, $\mu^{\text{MoC}}=\sum^{U}_{u=1}\mathds{1}_{u}\mu_{u}$, and $\epsilon^{\text{MoC}}\sim\text{Cauchy}(\textbf{0},\sum^{U}_{u=1}(\mathds{1}_{u}\gamma_{u}))$. Note that $\mathds{1}_{u}$ is a zero-one indicator and $\sum^{U}_{u=1}\mathds{1}_{u}=1$ and $\text{Prob}(\mathds{1}_{u}=1)=\pi_{u}$.
@@
From Eq. (5), we can observe that when $K$ is sufficiently large and $\overline{\alpha}_{K}$ correspondingly decreases to nearly zero, the distribution of $\hat{d}_{K}$ reaches the MoC distribution, i.e., $\hat{d}_{K}=\mu^{\text{MoC}}+\epsilon^{\text{MoC}}\sim\text{Cauchy}(\sum^{U}_{u=1}(\mathds{1}_{u}\mu_{u}),\sum^{U}_{u=1}(\mathds{1}_{u}\gamma_{u}))$. After the above MoC-based forward process, we can use the generated $\{\hat{D}_{k}\}^{K-1}_{k=1}$ as supervision signals to train the diffusion model $M_{\text{diff}}$ to learn the reverse process. More details about Eq. (5) can be found in Supplementary material. Such a forward process is only conducted to generate supervision signals for training the diffusion model, while we only need to conduct the reverse process during testing.
@@
However, it can still be difficult to generate $D_{0}$ by directly performing the reverse process from $\hat{D}_{K}$, because the object appearance features are lacking in $\hat{D}_{K}$. Such features can help constrain the model reverse process based on the input image to get accurate predictions. Thus we further leverage the appearance features from the image as context to guide $M_{\text{diff}}$ in the reverse process. Specifically, we reuse the features extracted from the keypoints distribution initializer as the appearance features $f_{\text{app}}$ and feed $f_{\text{app}}$ into the diffusion model, as shown in Fig. 3.
Our reverse process aims to generate a determinate distribution $D_{0}$ from the indeterminate distribution $\hat{D}_{K}$ (during training) or $D_{K}$ (during testing). Below we describe the reverse process during testing. We first obtain $f_{\text{app}}$ from the input image. Then to help the diffusion model to learn to perform denoising at each reverse step, following [18, 52], we generate the unique step embedding $f^{k}_{D}$ to inject the step number ($k$) information into the model. In this way, given a set of noisy keypoints coordinates $d_{k}\in\mathbb{R}^{N\times 2}$ drawn from ${D}_{k}$ at the $k^{th}$ step, we use diffusion model $M_{\text{diff}}$, conditioned on the step embedding $f^{k}_{D}$ and the object appearance features $f_{\text{app}}$, to recover ${d}_{k-1}$ from ${d}_{k}$ as:
${d}_{k-1}=M_{\text{diff}}({d}_{k},f_{\text{app}},f^{k}_{D})$ (6)